C&A sells 600 bottles of a dietary supplement per week at $100 per bottle. The supplement is ordered from a supplier who charges C&A $30 per order and $50 per bottle. C&A’s annual holding cost percentage is 40%. Assume C&A operates 50 weeks in a year. What is C&A’s total ordering and holding cost per year if C&A orders 500 bottles at a time?
Answer #1

Annual Requirement = Weekly Demand x Number of operational weeks

= 600 x 50 = 30,000 Bottles

Number of orders = Annual demand / Order Size

= 30,000 / 500

= 60

Ordering Cost = Number of orders x Ordering cost per order

= 60 x $30

= $1,800

Average Inventory = Lot Size / 2 = 500 / 2 = 250 Bottles

Holding cost per unit = Purchase Price x Holding cost percentage

= $50 x 40% = $20 per unit per annum

Holding cost of average inventory = Average Inventory x Holding cost per unit per annum

= 250 x $20 = $5,000

So, total ordering and holding cost per year when lot size is 500

= $1,800 + $5,000

= $6,800